# Exclusively Penalized Q-learning for Offline Reinforcement Learning

### v1

@@ 1 Introduction
* indicates equal contribution and $\dagger$ indicates the corresponding author: Seungyul Han.
Reinforcement learning (RL) is gaining significant attention for solving complex Markov decision process (MDP) tasks. Traditionally, online RL develops advanced decision-making strategies through continuous interaction with environments [1, 2, 3, 4, 5, 6]. However, in real-world scenarios, interacting with the environment can be costly, particularly in high-risk environments like disaster situations, where obtaining sufficient data for learning is challenging [7, 8]. This highlights the significance of research on offline setups, where policies are learned using pre-collected data without any direct interaction with the environment [9, 10]. In offline setups, policy actions not present in the data may introduce extrapolation errors, disrupting accurate value estimation by causing a large overestimation error in the value function, known as the distributional shift problem [11].
@@ 2.2 Distributional Shift Problem in Offline RL
In online RL, the optimal policy that maximizes $J(\pi)$ is found through iterative policy evaluation and policy improvement [4, 5]. For policy evaluation, the action value function is defined as $Q^{\pi}(s_{t},a_{t}):=\mathbb{E}_{s_{t},a_{t},s_{t+1},\cdots\sim\pi}[\sum^{\infty}_{l=t}\gamma^{l-t}R(s_{l},a_{l})|s_{t},~{}a_{t}]$. $Q^{\pi}$ can be estimated by iteratively applying the Bellman operator $\mathcal{B}^{\pi}$ to an arbitrary $Q$-function, where $(\mathcal{B}^{\pi}Q)(s,a):=R(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a),~{}a^{\prime}\sim\pi(\cdot|s^{\prime})}[Q(s^{\prime},a^{\prime})]$. The $Q$-function is updated to minimize the Bellman error using the dataset $D$, given by $\mathbb{E}_{s,a,s^{\prime}\sim D}\left[\left(Q(s,a)-\mathcal{B}^{\pi}Q(s,a)\right)^{2}\right]$. In offline RL, samples are generated by the behavior policy $\beta$ only, resulting in estimation errors in the $Q$-function for policy actions not present in the dataset $D$. The policy $\pi$ is updated to maximize the $Q$-function, incorporating the estimation error in the policy improvement step. This process accumulates positive bias in the $Q$-function as iterations progress [11].
@@ 3.1 Motivation: Necessity of Mitigating Unnecessary Estimation Bias
To visually demonstrate this, Fig. 1 depicts histograms of the fixed policy $\pi$ and the estimated behavior policy $\hat{\beta}$ for various $\pi$ and $\beta$ at the initial state $s_{0}$ on the Pendulum task with a single-dimensional action space in OpenAI Gym tasks [18], as cases (a), (b), and (c), along with the estimation bias in the $Q$-function for CQL with various penalizing factors $\alpha$. In this example, for all states except the initial state, we consider $\pi=\beta=\text{Unif}(-2,2)$. In each case, CQL only updates the $Q$-function with its penalty to evaluate $\pi$ in an offline setup, as shown in equation (1), and we plot the estimation bias of CQL, which represents the average difference between the learned $Q$-function and the actual average return $G_{0}$.
@@
From the results in Fig. 1, we observe that CQL suffers from unnecessary estimation bias in the $Q$-function for cases (a) and (b). In both cases, the histograms illustrate that policy actions are fully contained in the dataset $\hat{\beta}$, suggesting that the estimation error in the Bellman update is unlikely to occur even without any penalty. However, CQL introduces a substantial negative bias for actions near 00 where $\pi(0|s_{0})>\hat{\beta}(0|s_{0})$ and a positive bias for other actions. Furthermore, the bias intensifies as the penalty level $\alpha$ increases. In order to mitigate this bias, reducing the penalty level $\alpha$ to zero may seem intuitive in cases like Fig. 1(a) and (b). However, such an approach would be inadequate in cases like Fig. 1(c). In this case, because policy actions close to 0 are rare in the dataset, penalization is necessary to address overestimation caused by estimation errors in offline learning. Furthermore, this problem may become more severe in actual offline learning situations, as the policy continues to change as learning progresses, compared to situations where a fixed policy is assumed.
@@ 3.2 Exclusively Penalized Q-learning
To address the issue outlined in Section 3.1, our goal is to selectively give a penalty to the $Q$-function in cases like Fig. 1(c), where policy actions are insufficient in the dataset while minimizing unnecessary bias due to the penalty in scenarios like Fig. 1(a) and (b), where policy actions are sufficient in the dataset. To achieve this goal, we introduce a novel exclusive penalty $\mathcal{P}_{\tau}$ defined by
@@
$\mathcal{P}_{\tau}:=\underbrace{f_{\tau}^{\pi,\hat{\beta}}(s)}_{\textrm{penalty adaptation factor}}\cdot\underbrace{\left(\frac{\pi(a|s)}{\hat{\beta}(a|s)}-1\right)}_{\textrm{positive penalty if }\pi(a|s)>\hat{\beta}(a|s),~{}\textrm{otherwise negative}},$ (2)
Figure 2: An illustration of our methods: (a) Exclusive penalty on states where policy actions are lacking in the dataset (b) Prioritized dataset to prevent penalty increase due to policy concentration
@@
where $f_{\tau}^{\pi,\hat{\beta}}(s)=\mathbb{E}_{a\sim\pi(\cdot|s)}[\min(1.0,\exp(-(\log\hat{\beta}(a|s)-\tau)))]$ is a penalty adaptation factor that determines how much the policy actions are present in the dataset based on a threshold $\tau$. If the probability of estimated behavior policy $\hat{\beta}$ for policy actions exceeds the threshold $\tau$, i.e., policy actions are sufficiently present in the dataset, then $f_{\tau}^{\pi,\hat{\beta}}(s)$ will be smaller than 1 and reduce the amount of penalty as much as the amount by which $\hat{\beta}$ exceeds the threshold $\tau$ to avoid unnecessary bias introduced in Section 3.1. Otherwise, it will be 1 due to $\min(1.0,\cdot)$ to maintain the penalty since policy actions are insufficient in the dataset. The latter penalty term $\frac{\pi(a|s)}{\hat{\beta}(a|s)}-1$ imposes a positive penalty on the $Q$-function when $\pi(a|s)>\hat{\beta}(a|s)$, and otherwise, it increases the $Q$-function since the penalty is negative, as the Q-penalization method considered in CQL [15].
To elaborate further on our proposed penalty, Fig. 2(a) depicts the behavior of the penalty adaptation factor $f_{\tau}^{\pi,\hat{\beta}}$ across three different positions of policies ($\pi_{1}$, $\pi_{2}$, and $\pi_{3}$) and one fixed behavior policy $\hat{\beta}$. Note that CQL will impose a large penalty in all cases since $\pi_{i}>\hat{\beta}$ in this setup, regardless of the presence of policy actions in the dataset. This leads to unnecessary bias, as discussed in Section 3.1. In contrast, the proposed penalty $\mathcal{P}_{\tau}$ can mitigate such unnecessary bias even if $\pi_{i}$ is greater than $\beta$, as the penalty adaptation factor $f_{\tau}^{\pi,\hat{\beta}}$ decreases the penalty by the shaded area in Fig. 2(a). This area quantifies the degree to which the behavior policy probability $\hat{\beta}$ for policy actions exceeds the threshold $\tau$, representing the amount of policy actions in the dataset.
@@
Furthermore, to explain the role of the threshold $\tau$ in the proposed penalty $\mathcal{P}_{\tau}$, we consider two thresholds, $\tau_{1}$ and $\tau_{2}$ in Fig. 2(a). Here, each $\tau_{1}$ and $\tau_{2}$ considers the case where the number of data samples is $N_{1}$ and $N_{2}$, respectively ($N_{1}<N_{2}$). In the case of $\tau_{1}$, $\mathcal{P}_{\tau_{1}}$ only reduces the penalty for $\pi_{3}$ as much as the blue shaded area where $\hat{\beta}$ exceeds the threshold $\tau_{1}$ in the support of $\pi_{3}$. In other words, $f_{\tau_{1}}^{\pi_{1},\hat{\beta}}=f_{\tau_{1}}^{\pi_{2},\hat{\beta}}=1,$ and $f_{\tau_{1}}^{\pi_{3},\hat{\beta}}<1$. On the other hand, as the number of data samples increases from $N_{1}$ to $N_{2}$, more actions generated by the behavior policy $\beta$ will be stored in the dataset, so policy actions are more likely to be in the dataset. In this case, the threshold should be lowered from $\tau_{1}$ to $\tau_{2}$. As a result, $\hat{\beta}$ exceeds the threshold $\tau_{2}$ in the support of all policies $\pi_{i}$, and $\mathcal{P}_{\tau_{2}}$ reduces the penalty as much as the green shared area where $\hat{\beta}$ exceeds the threshold $\tau_{2}$ in the support of $\pi_{i}$, i.e., $f_{\tau_{2}}^{\pi_{3},\hat{\beta}}<f_{\tau_{2}}^{\pi_{1},\hat{\beta}}<f_{\tau_{2}}^{\pi_{2},\hat{\beta}}<1$. Thus, even without knowing the exact number of data samples, the proposed penalty $\mathcal{P}_{\tau}$ allows adjusting the penalty level appropriately according to the given number of data samples based on the threshold $\tau$.
@@
In order to demonstrate the $Q$-function convergence behavior of the proposed EPQ in more detail, we revisit the previous Pendulum task in Fig. 1. Fig. 4 shows the histogram of $\hat{\beta}$ and the penalty adaptation factor $f_{\tau}^{\pi,\hat{\beta}}(s)$ for Gaussian policy $\pi=N(\mu,0.2)$, where $\mu$ varies from $-2$ to $2$, with varying $\beta$. In Fig. 4(a), $f_{\tau}^{\pi,\hat{\beta}}(s)$ should be less than 1 for any policy mean $\mu$ since all policy actions are sufficient in the dataset. In 4(b), $f_{\tau}^{\pi,\hat{\beta}}(s)$ is less than 1 only if the $\hat{\beta}$ probability near the policy mean $\mu$ is high, and otherwise, $f_{\tau}^{\pi,\hat{\beta}}(s)$ is 1, which indicates the lack of policy action in the dataset. Thus, the result shows that $f_{\tau}^{\pi,\hat{\beta}}(s)$ reflects our motivation in Section 3.1 well. Moreover, Fig. 4 compares the estimation bias curves of CQL and EPQ with $\alpha=10$ in the scenarios presented in Fig. 1. CQL exhibits unnecessary bias for situations in Fig. 4(a) and (b) where no penalty is needed, as discussed in Section 3.1. Conversely, our proposed method effectively reduces estimation bias in these cases while appropriately maintaining the penalty for the scenario in Fig. 4(c) where penalization is required. This experiment demonstrates the effectiveness of our proposed approach, and the subsequent numerical results in Section 4 will numerically show that our method significantly reduces estimation bias in offline learning, resulting in improved performance.
@@ 3.3 Prioritized Dataset
In Section 3.2, EPQ effectively controls the penalty in the scenarios depicted in Fig. 4. However, in cases where the policy is highly concentrated on one side, as shown in Fig. 4, the estimation bias may not be completely eliminated due to the latter penalty term $\frac{\pi}{\hat{\beta}}-1$ in $\mathcal{P}_{\tau}$, as $\pi$ significantly exceeds $\hat{\beta}$. This situation, detailed in Fig. 2(b), arises when there is a substantial difference in the $Q$-function values among data actions. The policy is updated to maximize the $Q$-function, and thus the policy shifts towards the data action with a larger $Q$, resulting in a more significant penalty as illustrated in Fig. 2(b) for CQL. To further alleviate the penalty to reduce unnecessary bias in this situation, instead of applying a penalty based on $\hat{\beta}$, we introduce a penalty based on the prioritized dataset (PD) $\hat{\beta}^{Q}\propto\hat{\beta}\exp(Q)$. $\beta^{Q}$ prioritizes data actions with higher $Q$-values within the support of $\hat{\beta}$, as depicted in Fig. 2(b). Consequently, we anticipate that the penalty will be significantly mitigated, as the difference between $\pi$ and $\hat{\beta}^{Q}$ is much smaller than the difference between $\pi$ and $\hat{\beta}$. Following this intuition, we modify our penalty using PD as $\mathcal{P}_{\tau,~{}PD}:=f_{\tau}^{\pi,\hat{\beta}}(s)\cdot\left(\frac{\pi(a|s)}{\hat{\beta}^{Q}(a|s)}-1\right)$. It is important to note that the penalty adaptation factor $f_{\tau}^{\pi,\hat{\beta}}(s)$ remains unchanged since we use all data samples in the dataset for $Q$ updates. Additionally, we consider the prioritized dataset for the Bellman update to focus more on data actions with higher $Q$-function values for better performance as considered in [19]. Then, we can derive the final $Q$-loss function of EPQ with PD as

### v2

@@ 1 Introduction
* indicates equal contribution and $\dagger$ indicates the corresponding author: Seungyul Han.Special thanks to Whiyoung Jung from LG AI Research for providing experimental data used in this work.
Reinforcement learning (RL) is gaining significant attention for solving complex Markov decision process (MDP) tasks. Traditionally, online RL develops advanced decision-making strategies through continuous interaction with environments [1, 2, 3, 4, 5, 6]. However, in real-world scenarios, interacting with the environment can be costly, particularly in high-risk environments like disaster situations, where obtaining sufficient data for learning is challenging [7, 8]. In such setups, the need for exploration [9, 10, 11, 12] to discover optimal strategies often incurs additional costs, as agents must try various actions, some of which may be inefficient or risky [13, 14]. This highlights the significance of research on offline setups, where policies are learned using pre-collected data without any direct interaction with the environment [15, 16]. In offline setups, policy actions not present in the data may introduce extrapolation errors, disrupting accurate value estimation by causing a large overestimation error in the value function, known as the distributional shift problem [17].
@@ 2.2 Distributional Shift Problem in Offline RL
In online RL, the optimal policy that maximizes $J(\pi)$ is found through iterative policy evaluation and policy improvement [2, 3]. For policy evaluation, the action value function is defined as $Q^{\pi}(s_{t},a_{t}):=\mathbb{E}_{s_{t},a_{t},s_{t+1},\cdots\sim\pi}[\sum^{\infty}_{l=t}\gamma^{l-t}R(s_{l},a_{l})|s_{t},~{}a_{t}]$. $Q^{\pi}$ can be estimated by iteratively applying the Bellman operator $\mathcal{B}^{\pi}$ to an arbitrary $Q$-function, where $(\mathcal{B}^{\pi}Q)(s,a):=R(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a),~{}a^{\prime}\sim\pi(\cdot|s^{\prime})}[Q(s^{\prime},a^{\prime})]$. The $Q$-function is updated to minimize the Bellman error using the dataset $D$, given by $\mathbb{E}_{s,a\sim D}\left[\left(Q(s,a)-\mathcal{B}^{\pi}Q(s,a)\right)^{2}\right]$. In offline RL, samples are generated by the behavior policy $\beta$ only, resulting in estimation errors in the $Q$-function for policy actions not present in the dataset $D$. The policy $\pi$ is updated to maximize the $Q$-function, incorporating the estimation error in the policy improvement step. This process accumulates positive bias in the $Q$-function as iterations progress [17].
@@ 3.1 Motivation: Necessity of Mitigating Unnecessary Estimation Bias
To visually demonstrate this, Fig. 1 depicts histograms of the fixed policy $\pi$ and the estimated behavior policy $\hat{\beta}$ for various $\pi$ and $\beta$ at the initial state $s_{0}$ on the Pendulum task with a single-dimensional action space in OpenAI Gym tasks [24], as cases (a), (b), and (c), along with the estimation bias in the $Q$-function for CQL with various penalizing factors $\alpha$. In this example, for all states except the initial state, we consider $\pi=\beta=\text{Unif}(-2,2)$. In each case, CQL only updates the $Q$-function with its penalty to evaluate $\pi$ in an offline setup, as shown in equation (1), and we plot the estimation bias of CQL, which represents the average difference between the learned $Q$-function and the expected return $G_{0}$.
@@
From the results in Fig. 1, we observe that CQL suffers from unnecessary estimation bias in the $Q$-function for cases (a) and (b). In both cases, the histograms illustrate that policy actions are fully contained in the dataset $\hat{\beta}$, suggesting that the estimation error in the Bellman update is unlikely to occur even without any penalty. However, CQL introduces a substantial negative bias for actions near 00 where $\pi(0|s_{0})>\hat{\beta}(0|s_{0})$ and a positive bias for other actions. Furthermore, the bias intensifies as the penalty level $\alpha$ increases. In order to mitigate this bias, reducing the penalty level $\alpha$ to zero may seem intuitive in cases like Fig. 1(a) and Fig. 1(b). However, such an approach would be inadequate in cases like Fig. 1(c). In this case, because policy actions close to 0 are rare in the dataset, penalization is necessary to address overestimation caused by estimation errors in offline learning. Furthermore, this problem may become more severe in actual offline learning situations, as the policy continues to change as learning progresses, compared to situations where a fixed policy is assumed.
@@ 3.2 Exclusively Penalized Q-learning
To address the issue outlined in Section 3.1, our goal is to selectively give a penalty to the $Q$-function in cases like Fig. 1(c), where policy actions are insufficient in the dataset while minimizing unnecessary bias due to the penalty in scenarios like Fig. 1(a) and Fig. 1(b), where policy actions are sufficient in the dataset. To achieve this goal, we introduce a novel exclusive penalty $\mathcal{P}_{\tau}$ defined by
@@
$\mathcal{P}_{\tau}:=\underbrace{f_{\tau}^{\pi,\hat{\beta}}(s)}_{\textrm{penalty adaptation factor}}\cdot\underbrace{\left(\frac{\pi(a|s)}{\hat{\beta}(a|s)}-1\right)}_{\textrm{penalty term}},$ (2)
Figure 2: An illustration of our exclusive penalty: (a) The log-probability of $\hat{\beta}$ and the thresholds $\tau_{1}$ and $\tau_{2}$ according to the number of data samples $N_{1}$ and $N_{2}$, where $N_{1}<<N_{2}$. (b) The penalty adaptation factor $f^{\pi,\hat{\beta}}_{\tau}$ which represents the amount of adaptive penalty, indicating how much $\log\hat{\beta}$ exceeds the threshold $\tau$. Three different policies $\pi_{i},~{}i=1,2,3$, are considered.
@@
where $f_{\tau}^{\pi,\hat{\beta}}(s)=\mathbb{E}_{a\sim\pi(\cdot|s)}[x_{\tau}^{\hat{\beta}}]$ is a penalty adaptation factor for a given $\hat{\beta}$ and policy $\pi$. Here, $x_{\tau}^{\hat{\beta}}=\min(1.0,\exp(-(\log\hat{\beta}(a|s)-\tau)))$ represents the amount of adaptive penalty that is reduced as log $\hat{\beta}$ exceeds the threshold $\tau$. Thus, the adaptation factor $f^{\pi,\hat{\beta}}_{\tau}$ indicates the average penalty that policy actions should receive. If the probability of estimated behavior policy $\hat{\beta}$ for policy actions exceeds the threshold $\tau$, i.e., policy actions are sufficiently present in the dataset, then $x_{\tau}^{\hat{\beta}}$ will be smaller than 1 and reduce the amount of penalty as much as the amount by which $\hat{\beta}$ exceeds the threshold $\tau$ to avoid unnecessary bias introduced in Section 3.1. Otherwise, it will be 1 due to $\min(1.0,\cdot)$ to maintain the penalty since policy actions are insufficient in the dataset. The latter penalty term $\frac{\pi(a|s)}{\hat{\beta}(a|s)}-1$, positive if $\pi(a|s)>\hat{\beta}(a|s)$ and otherwise negative, imposes a positive penalty on the $Q$-function when $\pi(a|s)>\hat{\beta}(a|s)$, and otherwise, it increases the $Q$-function since the penalty is negative, as the Q-penalization method considered in CQL [21].
To elaborate further on our proposed penalty, Fig. 2(a) depicts the log-probability of $\hat{\beta}$ and the thresholds $\tau$ used for penalty adaptation, with $N$ representing the number of data points. In Fig. 2(a), if the log-probability $\log\hat{\beta}$ of an action $a\in\mathcal{A}$ exceeds the threshold $\tau$, this indicates that the action $a$ is sufficiently represented in the dataset. Thus, we reduce the penalty for such actions. Furthermore, as shown in Fig. 2(a), when the number of actions increase from $N_{1}$ to $N_{2}$, the threshold for determining "enough data" decreases from $\tau_{1}$ to $\tau_{2}$, even if the data distribution remains unchanged.
@@
Furthermore, to explain the role of the threshold $\tau$ in the proposed penalty $\mathcal{P}_{\tau}$, we consider two thresholds, $\tau_{1}$ and $\tau_{2}$. In Fig. 2(b), which illustrates the proposed penalty adaptation factor $f^{\pi,\hat{\beta}}_{\tau_{1}}$ and $f^{\pi,\hat{\beta}}_{\tau_{2}}$ for thresholds $\tau_{1}$ and $\tau_{2}$, $x_{\tau_{1}}^{\hat{\beta}}$ is larger than $x_{\tau_{2}}^{\hat{\beta}}$ because $\tau_{1}>\tau_{2}$. As a result, in the case of $\tau_{1}$, $\mathcal{P}_{\tau_{1}}$ only reduces the penalty for $\pi_{3}$. In other words, $f_{\tau_{1}}^{\pi_{1},\hat{\beta}}=f_{\tau_{1}}^{\pi_{2},\hat{\beta}}=1,$ and $f_{\tau_{1}}^{\pi_{3},\hat{\beta}}<1$. On the other hand, as the number of data samples increases from $N_{1}$ to $N_{2}$, more actions generated by the behavior policy $\beta$ will be stored in the dataset, so policy actions are more likely to be in the dataset. In this case, the threshold should be lowered from $\tau_{1}$ to $\tau_{2}$. As a result, $\hat{\beta}$ exceeds the threshold $\tau_{2}$ in the support of all policies $\pi_{i}$, and $\mathcal{P}_{\tau_{2}}$ reduces the penalty in the support of all policies $\pi_{i}$, i.e., $f_{\tau_{2}}^{\pi_{3},\hat{\beta}}<f_{\tau_{2}}^{\pi_{1},\hat{\beta}}<f_{\tau_{2}}^{\pi_{2},\hat{\beta}}<1$. Thus, even without knowing the exact number of data samples, the proposed penalty $\mathcal{P}_{\tau}$ allows adjusting the penalty level appropriately according to the given number of data samples based on the threshold $\tau$.
@@
In order to demonstrate the $Q$-function convergence behavior of the proposed EPQ in more detail, we revisit the previous Pendulum task in Fig. 1. Fig. 4 shows the histogram of $\hat{\beta}$ and the penalty adaptation factor $f_{\tau}^{\pi,\hat{\beta}}(s)$ for Gaussian policy $\pi=N(\mu,0.2)$, where $\mu$ varies from $-2$ to $2$, with varying $\beta$. In Fig. 4(a), $f_{\tau}^{\pi,\hat{\beta}}(s)$ should be less than 1 for any policy mean $\mu$ since all policy actions are sufficient in the dataset. In 4(b), $f_{\tau}^{\pi,\hat{\beta}}(s)$ is less than 1 only if the $\hat{\beta}$ probability near the policy mean $\mu$ is high, and otherwise, $f_{\tau}^{\pi,\hat{\beta}}(s)$ is 1, which indicates the lack of policy action in the dataset. Thus, the result shows that $f_{\tau}^{\pi,\hat{\beta}}(s)$ reflects our motivation in Section 3.1 well. Moreover, Fig. 4 compares the estimation bias curves of CQL and EPQ with $\alpha=10$ in the scenarios presented in Fig. 1. CQL exhibits unnecessary bias for situations in Fig. 4(a) and Fig. 4(b) where no penalty is needed, as discussed in Section 3.1. Conversely, our proposed method effectively reduces estimation bias in these cases while appropriately maintaining the penalty for the scenario in Fig. 4(c) where penalization is required. This experiment demonstrates the effectiveness of our proposed approach, and the subsequent numerical results in Section 4 will numerically show that our method significantly reduces estimation bias in offline learning, resulting in improved performance.
@@ 3.3 Prioritized Dataset
Figure 5: An illustration of the prioritized dataset. As the policy focuses on actions with maximum $Q$-values, the difference between $\hat{\beta}$ and $\pi$ becomes substantial, inducing large penalty: (a) The change of data distribution from $\hat{\beta}$ (w/o PD) to $\hat{\beta}^{Q}$ (with PD) (b) The corresponding penalty graphs for $\hat{\beta}$ (w/o PD) and $\hat{\beta}^{Q}$ (with PD).
In Section 3.2, EPQ effectively controls the penalty in the scenarios depicted in Fig. 4. However, in cases where the policy is highly concentrated on one side, as shown in Fig. 4, the estimation bias may not be completely eliminated due to the latter penalty term $\frac{\pi}{\hat{\beta}}-1$ in $\mathcal{P}_{\tau}$, as $\pi$ significantly exceeds $\hat{\beta}$. This situation, detailed in Fig. 5, arises when there is a substantial difference in the $Q$-function values among data actions. As the policy is updated to maximize the $Q$-function, the policy shifts towards the data action with a larger $Q$, resulting in a more significant penalty for CQL. To further alleviate the penalty to reduce unnecessary bias in this situation, instead of applying a penalty based on $\hat{\beta}$, we introduce a penalty based on the prioritized dataset (PD) $\hat{\beta}^{Q}\propto\hat{\beta}\exp(Q)$. As shown in Fig. 5(a), which illustrates the difference between the original data distribution $\hat{\beta}$ and the modified data distribution $\hat{\beta}^{Q}$ after applying PD, $\beta^{Q}$ prioritizes data actions with higher $Q$-values within the support of $\hat{\beta}$. According to Fig. 5(a), when the policy $\pi$ focuses on specific actions, the penalty $\frac{\pi}{\hat{\beta}}-1$ increases significantly, as depicted in Fig. 5(b). In contrast, by applying PD, $\hat{\beta}$ is adjusted to approach $\hat{\beta}^{Q}\propto\beta\exp(Q)$, aligning the data distribution more closely with the policy $\pi$. Consequently, we anticipate that the penalty will be significantly mitigated, as the difference between $\pi$ and $\hat{\beta}^{Q}$ is much smaller than the difference between $\pi$ and $\hat{\beta}$. Following this intuition, we modify our penalty using PD as $\mathcal{P}_{\tau,~{}PD}:=f_{\tau}^{\pi,\hat{\beta}}(s)\cdot\left(\frac{\pi(a|s)}{\hat{\beta}^{Q}(a|s)}-1\right)$. It is important to note that the penalty adaptation factor $f_{\tau}^{\pi,\hat{\beta}}(s)$ remains unchanged since we use all data samples in the dataset for $Q$ updates. Additionally, we consider the prioritized dataset for the Bellman update to focus more on data actions with higher $Q$-function values for better performance as considered in [25]. Then, we can derive the final $Q$-loss function of EPQ with PD as